OpenStudy (anonymous):

FAN AND MEDAL 3w-6=-1.5 how would I work this out

pooja195 (pooja195):

start by adding 6 to both sides

pooja195 (pooja195):

\[\huge~\rm~3w-6+6=-1.5+6\]

OpenStudy (anonymous):

Thanks you im already a fan but ill medal

pooja195 (pooja195):

I just want to make sure I help you :)

pooja195 (pooja195):

So when we add 6 to both sides what do we get?

OpenStudy (anonymous):

12

pooja195 (pooja195):

Not quite... -6+6=?

pooja195 (pooja195):

They cancel out... -1.5+6=?

rvc (rvc):

when u give me 6 pie and take six pie from me what do u give me?

OpenStudy (anonymous):

0

rvc (rvc):

similarily -6+6=0

pooja195 (pooja195):

ok -1.5+6=?

pooja195 (pooja195):

\[\huge~\rm~3w=-1.5+6=?\]

rvc (rvc):

same way u gave me 6 n took 3/2tth part away what do i get?

OpenStudy (anonymous):

3?

rvc (rvc):

6-1.5 6.0-1.5=?

pooja195 (pooja195):

Ill let @rvc take over :) goodluck .

rvc (rvc):

nope im busy @pooja195

pooja195 (pooja195):